# Vikhr: The Family of Open-Source Instruction-Tuned Large Language Models for Russian

### v2

@@ Abstract
There has been a surge in the development of various Large Language Models (LLMs). However, text generation for languages other than English often faces significant challenges, including poor generation quality and the reduced computational performance due to the disproportionate representation of tokens in model’s vocabulary. In this work, we address these issues and introduce Vikhr, a new state-of-the-art open-source instruction-tuned LLM designed specifically for the Russian language. “Vikhr” refers to the name of the Mistral LLM series and means “strong gust of wind.” Unlike previous efforts for Russian that utilize computationally inexpensive LoRA adapters on top of English-oriented models, Vikhr features an adapted tokenizer vocabulary and undergoes the continued pre-training and instruction tuning of all weights. This approach not only enhances the model’s performance but also significantly improves its computational and contextual efficiency. The remarkable performance of Vikhr across various Russian-language benchmarks can also be attributed to our efforts in expanding instruction datasets and corpora for continued pre-training. Vikhr not only sets the new state of the art among open-source LLMs for Russian, but even outperforms some proprietary closed-source models on certain benchmarks. The model weights, instruction sets, and code are publicly available [^1] https://huggingface.co/Vikhrmodels.
@@ 1 Introduction
Instruction tuning has unlocked in Large Language Models (LLMs) vast zero-shot capabilities without the need of careful prompt engineering Ouyang et al. (2022). The most rapid research and development efforts are currently devoted to English LLMs. There has been a surge in English open-source models: Llama series Touvron et al. (2023a, b), Mistral series Jiang et al. (2023), Vicuna series Chiang et al. (2023), etc. This growth is driven by the abundance of raw training data in English and dedicated efforts to create comprehensive sets of instruction-output pairs. Despite the fact that LLMs oriented on English have some multilingual capabilities Zhao et al. (2024) due to small portions of texts in various languages leaked into their training datasets Touvron et al. (2023a), their overall performance in these languages remains relatively low. Although they can usually generate portions of coherent texts, these models struggle with reasoning in non-English languages, lack culture-specific knowledge, and are highly inefficient in terms of tokenization. This inefficiency arises due to the way bite-pair tokenization algorithms work: they split the infrequent words into multiple tokens. Since multilingual data typically represents a small portion of the training dataset, non-English words are often split in many pieces. This leads to more steps during prompt processing and text generation, shorter effective context windows, and ultimately lower quality Tikhomirov and Chernyshev (2023); Petrov et al. (2024). This disparity places non-English languages at a disadvantage.
@@
There is a research direction focused on developing multilingual LLMs that work well for multiple popular languages: BLOOMz Muennighoff et al. (2023), mGPT Shliazhko et al. (2022), Bactrian-X Li et al. (2023), PALO Maaz et al. (2024), Aya101 from CohereAI Üstün et al. (2024), etc. These models are typically trained on rich multilingual datasets and are less skewed towards English. However, when aiming to perform well across multiple languages simultaneously, these models must still share their vocabulary and parameters. This often hinders their performance for each particular language in isolation, especially for the popular smaller model sizes, such as 7B and 13B.
@@
The goal of maximizing the LLM performance for a specific language within a certain number of parameters has led researchers to develop bi-lingual LLMs. For example, Jais Sengupta et al. (2023) focus only on English and Arabic. The inclusion of English data in pre-training alongside Arabic data is motivated by the significantly larger volume of English data available. This helps LLMs substantially enhance skills such as logical and common sense reasoning, which are also applied when generating text in Arabic.
@@
Russian is one of the high-resource languages and is typically represented in multilingual LLMs. Additionally, there are several proprietary closed-source LLMs, such as MTS AI, GigaChat, and YandexGPT, that meet or even surpass their English-oriented flagship competitors when it comes to text processing and generation in Russian. However, controllable research often requires white-box access to LLM logits and layer outputs, the ability to modify weights and a model architecture, and consistent answers for reproducibility, which is often impossible in closed-source LLMs due to their constant development and retirement. There are only a few open-source LLMs designed for Russian (such as Saiga Gusev (2023), ruGPT AI Forever (2022), ruadapt Tikhomirov and Chernyshev (2023)). Of these, only Saiga and ruadapt are instruction-tuned.
@@
This work aims to build an efficient and effective open-source instruction-following LLM for Russian facilitating multilingual natural language processing research. Building even a small LLM that targets a particular language from scratch requires a lot of computational resources. Consequently, many researchers simply fine-tune LoRA adapters Hu et al. (2021) for English-oriented LLMs on some language-specific data. While this approach can improve model generation quality, it does not address computational inefficiency because the tokenizer and model vocabulary remain unchanged. In contrast, our approach not only fine-tunes a base LLM on Russian language data but also reconstructs its underlying tokenizer and vocabulary, alongside suggesting an improved method for continued pre-training. Additionally, we have significantly expanded the available Russian datasets for instruction tuning. The developed LLM achieves state-of-the-art results for the Russian language among other open-source counterparts across a wide range of benchmarks.
@@
  • We have constructed Vikhr – a state-of-the-art open-source instruction-following LLM oriented on the Russian language. In addition to its high generation quality, Vikhr features an efficient tokenizer that enables rapid text generation and good context utilization.
  • We have developed a pipeline for adapting English-oriented LLMs to the Russian language. The pipeline implements vocabulary adaptation, continued pre-training with regularization to prevent “catastrophic forgetting”, and instruction tuning.
@@
  • We conducted an extensive evaluation of several open-source LLMs on evaluation benchmarks for Russian, demonstrating that Vikhr achieves new state-of-the-art results.
@@ 2 Related Work
One of the first notable series of generative LLMs for Russian is ruGPT AI Forever (2022); Zmitrovich et al. (2023). The authors created several models trained for the vanilla language modelling task with the sizes of up to 13b. The models were created from the scratch and trained on large Russian corpora. They are able to handle the linguistic nuances of Russian more effectively than multilingual models Muennighoff et al. (2022). Since the training data was mostly in Russian, these models have efficient tokenization, but the lack of multilingual data (e.g. in English) limits their performance. ruGPT models are not instruction tuned.
Gusev (2023) suggests to leverage reasoning capabilities of existing English-oriented LLMs and adapt them to the Russian language by training LoRA adapters. They also create an Alpaca-like set of Russian instruction-output pairs and performed instruction tuning. They have established the Saiga model series, which has a competitive performance and used to be a reasonable choice for off-the-shelf open-source Russian LLM for the past year. However, the tokenizer in theses models is not adapted, so they experience issues with context and computational efficiency.
Tikhomirov and Chernyshev (2023) address these issues in Saiga. In addition to model tuning on Russian data, they also adapt the model tokenizer. They note that improving tokenization helps to both improve the efficiency of the model and its performance while reducing memory consumption. However, during continued pre-training, the authors freeze the model weights except LM heads and token embeddings, which probably results in the suboptimal performance.
In this work, we take advantage of pre-trained English-oriented LLMs, adapt LLM tokenizer for better computational efficiency, leverage continued pre-training on vast Russian-language corpora with regularization for preventing “catastrophic forgetting”, construct a novel extended set of Russian instruction-output pairs, and perform instruction tuning. The created LLM adaptation pipeline along with the data for continued pre-training and instruction tuning enables Vikhr to achieve new state-of-the-art results for Russian, maintain high performance for English, and demonstrate high computational efficiency.
@@ 3 LLM Construction Pipeline
The construction of Vikhr starts from one of English-oriented LLMs. In this work, we discuss the Vikhr model based on Mistral 7B. The strong logical and common reasoning capabilities, as well as the extensive world knowledge present in these LLMs provide an excellent starting point for our model. These features partially transfer to Vikhr, enhancing its performance in generating text in Russian. The process of LLM adaptation to Russian starts with the vocabulary adaptation. Then we perform continued pre-training of the LLM on large Russian datasets to mitigate the vocabulary shift and introduce culture specific knowledge. Finally, we perform fine-tuning of Vikhr on a set of instruction-output pairs in Russian.
@@ 3.1 Vocabulary Adaptation
Original Sentence 31 Машинное обучение изменяет мир
Mistral Tokenizer 13 [‘Ма’, ‘шин’, ‘ное’, ‘об’, ‘у’, ‘чение’, ‘из’, ‘мен’, ‘я’, ‘ет’, ‘ми’, ‘р’ ]
@@
Table 1: Tokenizer comparisons between the original Mistral model and Vikhr
Figure 1: The Vikhr tokenizer efficiency in comparison to tokenizers of other models.
The big drawback of English-oriented LLMs is that each Russian word would be split into multiple tokens: a common case is when symbols in the word become an individual tokens (see example in Table 1). This slows down the generation by multiple times, reduces the amount of information that could be stored in the context, and drastically hurts the generation quality.
@@ 3.2 Continued Pre-training
The new vocabulary requires also new embedding matrices and LM heads. The tokens that were present in the original vocabulary are initialized with the old embeddings, the new tokens are initialized by averaging the embeddings of their pieces in the original embedding matrix Hewitt (2021). The similar approach is also applied to LM heads. Training model with these modifications requires much more computational resources than the mainstream technique for adaptation of LLMs to new languages based on LoRA adapters Hu et al. (2021), as it requires to perform continued pre-training of the whole model and on much more language-specific data to mitigate the shift in the vocabulary.
The dataset for continued pre-training is constructed from Russian Wikipedia, news articles, scientific papers, top 100k up-voted posts on Habr, and some other sources. The statistics of these datasets is presented in Table 2. The total number of tokens used for this step is 11 billion.
We note that the continued pre-training of a LLM might partially eliminate the reasoning capabilities present in the original English-oriented model. This drastically affects the model performance. In our preliminary experiments, continued pre-training may result even in worse performance on Russian benchmarks compared to the original model. To alleviate the “catastrophic forgetting”, we use the loss regularization with KL penalty between the probability distribution of Vikhr and the reference English-oriented original LLM:
@@
In practice, we implement this approach using the SLERP interpolation of model losses Goddard et al. (2024).
@@
To speed up the process of continued pre-training, we use an optimized Flash attention implementation [^2] https://huggingface.co/docs/optimum/bettertransformer/tutorials/convert. As an optimization algorithm, we leverage AdamW as it trades some memory efficiency in favor of robustness to the hyperparameter choice. The hyperparameters used for continued pre-training are presented in Table 3.
@@ 3.3 Instruction Tuning
Instruction tuning is an essential step in reaching high zero-shot performance with LLMs. It also allows to obtain more natural communication with the model without complex prompting. Further fine-tuning techniques such as RLHF Ouyang et al. (2022), which require input from the assessors, are also crucial for such tasks as multicriteria alignment. However, the most significant performance gains are still achieved through instruction tuning Jha et al. (2023).
Previously, Gusev (2023) constructed an open-source set of instruction-output pairs for the Russian language (Saiga). The core Saiga dataset was created similar to Alpaca by querying ChatGPT (gpt-3.5-turbo) Taori et al. (2023). In this work, we extend this set by translating two English instruction datasets. First, we translated instructions for the FLAN model Wei et al. (2021) and generated answers in Russian using ChatGPT. Originally, FLAN instructions were constructed automatically from annotated datasets using templates to facilitate multitask and zero-shot capabilities of seq2seq models. Later, it was shown that this data also helps to improve decoder-only chat-oriented models as well. Second, we construct Veles [^3] https://huggingface.co/datasets/Vikhrmodels/Veles-2.5 by translating the English OpenHermes Teknium (2023) instruction dataset. We also include without translation Nectar [^4] https://huggingface.co/datasets/berkeley-nest/Nectar Zhu et al. (2023) – the English instruction dataset. It helps to keep the performance of Vikhr high also for English. Since the majority of the outputs were machine generated there are many low quality outputs. To mitigate this problem, we filtered out low quality pairs using a reward model trained on human data. The statistics of the Vikhr instruction datasets is presented in Table 4.
@@
Table 4: The statistics of instruction datasets.
@@
Vikhr-7B-instruct_0.2 true pt+sft 0.62 0.80 0.231 0.67 0.485

### v4

# Vikhr: Constructing a State-of-the-art Bilingual Open-Source Instruction-Following Large Language Model for Russian

@@ Abstract
There has been a surge in the development of various Large Language Models (LLMs). However, text generation for languages other than English often faces significant challenges, including poor generation quality and reduced computational performance due to the disproportionate representation of tokens in the model’s vocabulary. In this work, we address these issues by developing a pipeline for adaptation of English-oriented pre-trained models to other languages and constructing efficient bilingual LLMs. Using this pipeline, we construct Vikhr, a state-of-the-art bilingual open-source instruction-following LLM designed specifically for the Russian language. “Vikhr” refers to the name of the Mistral LLM series and means a “strong gust of wind.” Unlike previous Russian-language models that typically rely on LoRA adapters on top of English-oriented models, sacrificing performance for lower training costs, Vikhr features an adapted tokenizer vocabulary and undergoes the continued pre-training and instruction tuning of all weights. This not only enhances the model’s performance but also significantly improves its computational and contextual efficiency. The remarkable performance of Vikhr across various Russian-language benchmarks can also be attributed to our efforts in expanding instruction datasets and corpora for continued pre-training. Vikhr not only sets the new state of the art among open-source LLMs for Russian but even outperforms some proprietary closed-source models on certain benchmarks. The model weights, instruction sets, and code are publicly available [^1] https://huggingface.co/Vikhrmodels.
footnotetext: $\diamondsuit$ Equal contribution
@@ 1 Introduction
Instruction tuning has unlocked in Large Language Models (LLMs) vast zero-shot capabilities without the need for careful prompt engineering Ouyang et al. (2022). The most rapid research and development efforts are currently devoted to English LLMs. There has been a surge in English open-source models: Llama series Touvron et al. (2023a, b), Mistral series Jiang et al. (2023), Vicuna series Chiang et al. (2023), etc. This growth is driven by the abundance of raw training data in English and dedicated efforts to create extensive sets of instruction-output pairs. Even though LLMs oriented on English have some multilingual capabilities Zhao et al. (2024) due to the presence of small amounts of text in various languages within their training datasets Touvron et al. (2023a), their overall performance in non-English languages remains relatively limited. Although they can usually generate portions of coherent texts, these models struggle with reasoning in non-English languages, lack culture-specific knowledge, and are highly inefficient in terms of tokenization. This inefficiency stems from how byte-pair tokenization algorithms operate, as they break down infrequent words into multiple tokens. Since multilingual data typically represents a small portion of the training dataset, non-English words are often split into many pieces. As a result, this increases the number of steps during prompt processing and text generation, reduces the effective context window, and ultimately degrades overall performance Tikhomirov and Chernyshev (2023); Petrov et al. (2024). This disparity places non-English languages at a disadvantage.
@@
There is also a research direction focused on developing multilingual LLMs designed to perform well across multiple popular languages: BLOOMz Muennighoff et al. (2023), mGPT Shliazhko et al. (2022), Bactrian-X Li et al. (2023), PALO Maaz et al. (2024), Aya101 from CohereAI Üstün et al. (2024), etc. These models are typically trained on rich multilingual datasets and are less skewed towards English. However, when aiming to perform well across multiple languages simultaneously, these models must still share their vocabulary and parameters. This often hinders their performance for each particular language in isolation, especially for the popular smaller model sizes, such as 7B and 13B.
@@
The aim of maximizing the LLM performance for a specific language within a certain number of parameters has led researchers to develop bilingual LLMs Sengupta et al. (2023); Pieri et al. (2024); Faysse et al. (2024). These LLMs prioritize a regional language, e.g. JaisSengupta et al. (2023) focuses on Arabic, but they are trained also on English data. The inclusion of English data in pre-training alongside regional language data is motivated by the significantly larger volume of English data available. This helps LLMs substantially enhance skills such as logical and common sense reasoning, which are also applied when generating text in a regional language. Bilingual LLMs is a perspective direction as they can remain small and efficient, but at the same time comprehensively capture linguistic nuances and cultural contexts of the regional language.
This work seeks to develop a pipeline for adapting English LLMs to other languages facilitating the development of bilingual LLMs. Specifically, we aim to build an instruction-following bilingual LLM for Russian and English that could be used for multilingual natural language processing research.
@@
Russian is one of the high-resource languages and is typically represented in multilingual LLMs. Additionally, there are several proprietary closed-source LLMs, such as MTS AI, GigaChat, and YandexGPT, that meet or even surpass their English-oriented flagship competitors when it comes to text processing and generation in Russian. However, controllable research often requires white-box access to LLM logits and layer outputs, the ability to modify weights and model architecture, and consistent answers for reproducibility, which is often impossible in closed-source LLMs due to their constant development and retirement. There are only a few open-source LLMs designed for Russian: Saiga Gusev (2023), ruGPT AI Forever (2022), ruadapt Tikhomirov and Chernyshev (2023), and some others. Of these, only Saiga and ruadapt are instruction-tuned. We aim to fill the lack of instruction-tuned open-source LLM for Russian that is both efficient and effective.
@@
Building even a small LLM tailored to a specific language from scratch demands a lot of computational resources. Consequently, many researchers opt to fine-tune LoRA adapters Hu et al. (2021) for English-oriented LLMs using some language-specific data. While this approach can improve model generation quality, it does not address computational inefficiency because the tokenizer and model vocabulary remain unchanged. In contrast, our approach not only fine-tunes a base LLM on Russian language data but also reconstructs its underlying tokenizer and vocabulary, alongside suggesting an improved method for continued pre-training. Additionally, we have significantly expanded the available Russian datasets for instruction tuning. The developed LLM achieves state-of-the-art results for the Russian language among other open-source counterparts across a wide range of benchmarks.
@@
  • We have developed a pipeline for adapting English-oriented LLMs to other languages. The pipeline implements vocabulary adaptation, continued pre-training with regularization to prevent “catastrophic forgetting”, and instruction tuning.
  • Using the pipeline, we have constructed Vikhr – a state-of-the-art open-source instruction-following LLM oriented on the Russian language. In addition to its high generation quality, Vikhr features an efficient tokenizer that enables rapid text generation and good context utilization.
@@
  • We have constructed two evaluation benchmarks for Russian LLMs by translating the English MMLU Hendrycks et al. (2020) and MMLU-pro Wang et al. (2024b) benchmarks.
  • We conducted an extensive evaluation of several open-source LLMs on evaluation benchmarks for Russian, demonstrating that Vikhr achieves new state-of-the-art results. Ablation studies confirm the effectiveness and validity of the individual components within our LLM adaptation pipeline.
@@ 2 Related Work
One of the first prominent series of generative LLMs for Russian is ruGPT AI Forever (2022); Zmitrovich et al. (2023). The authors developed several models, trained on the standard language modeling task, with sizes reaching up to 13 billion parameters. These models were created from the scratch and trained on a large Russian corpus, enabling them to capture the linguistic nuances of Russian more effectively than multilingual models. Additionally, since the training data was mostly in Russian, these models also have efficient tokenization. However, the lack of multilingual data (e.g. in English) limits their performance. Notably, the ruGPT models are not instruction-tuned.
Gusev (2023) suggests to leverage reasoning capabilities of existing English-oriented LLMs and adapt them to the Russian language by training LoRA adapters. They created an Alpaca-like set of Russian instruction-output pairs and performed instruction tuning on it. As a result, they have established the series of models called Saiga, which demonstrate the competitive performance and used to be a reasonable choice for an off-the-shelf open-source Russian LLMs for the past years. However, the tokenizer in these models is not adapted, so they experience issues with the context size and computational efficiency.
Tikhomirov and Chernyshev (2023) address these issues in Saiga. In addition to model tuning on Russian data, they also adapted the model tokenizer. They note that improving tokenization helps to both enhance the efficiency of the model and its performance while reducing memory consumption. However, during continued pre-training, the authors freezed the model weights, except for LM heads and token embeddings, which leads to the suboptimal performance.
In this work, we take advantage of pre-trained English-oriented LLMs, adapt LLM tokenizer for better computational and contextual efficiency, leverage continued pre-training on vast Russian-language corpora with regularization for preventing “catastrophic forgetting”, construct a novel extended set of Russian instruction-output pairs, and perform instruction tuning. The created LLM adaptation pipeline along with the data for continued pre-training and instruction tuning enables Vikhr to achieve new state-of-the-art results for Russian, maintain high performance for English, and demonstrate high computational efficiency.
@@ 3 LLM Construction Pipeline
The construction of Vikhr starts from one of the English-oriented LLMs. In this work, we discuss the Vikhr model based on Mistral 7B Jiang et al. (2023). The strong logical and common-sense reasoning capabilities, as well as the extensive world knowledge present in Mistral LLMs provide an excellent starting point for our model. These features partially transfer to Vikhr, enhancing its performance in generating text in Russian. The process of the LLM adaptation to Russian starts with the vocabulary adaptation. Then we perform continued pre-training of the LLM on large Russian datasets to mitigate the vocabulary shift and introduce culture specific knowledge. Finally, we perform fine-tuning of Vikhr on a set of instruction-output pairs in Russian.
@@ 3.1 Vocabulary Adaptation
Original Sentence 31 (ru) Машинное обучение изменяет мир
[(en) Machine learning changes the world.
Mistral Tokenizer 13 [‘Ма’, ‘шин’, ‘ное’, ‘об’, ‘у’, ‘чение’, ‘из’, ‘мен’, ‘я’, ‘ет’, ‘ми’, i ‘р’ ]
@@
Table 1: A comparison of tokenization between the original Mistral model and Vikhr.
Figure 1: The efficiency of the Vikhr tokenizer for Russian in comparison to tokenizers of other models.
The big drawback of English-oriented LLMs is that each Russian word would be split into multiple tokens: a common case is when symbols in the word become individual tokens (see example in Table 1). This slows down the generation by multiple times, reduces the amount of information that could be stored in the context, and drastically hurts the generation quality.
@@ 3.2 Continued Pre-training
The new vocabulary requires new embedding matrices and LM heads. The tokens that were present in the original vocabulary are initialized with the old embeddings, the new tokens are initialized by averaging the embeddings of their pieces in the original embedding matrix Hewitt (2021). The similar approach is also applied to LM heads. Training models with these modifications demands much more computational resources than the mainstream method of adapting LLMs to new languages using LoRA adapters Hu et al. (2021). This is because it involves continued pre-training of the entire model and requires much more language-specific data to mitigate the shift in the vocabulary.
The dataset for continued pre-training is constructed from high-quality sources, including Russian Wikipedia, news articles, scientific papers from peer-reviewed journals and conferences, and top 100k up-voted posts on Habr – a popular online blog community focused on technology, software development, and science. The statistics of these datasets is presented in Table 2. We performed deduplication of the collection on the level of paragraphs using the MIHash algorithm Cakir et al. (2017). Furthermore, we performed filtration of the collected data. For this purpose, we annotated 20k documents using GPT-4-turbo with the aim to assess their informativeness, usefulness for studying, grammatical correctness, style, and safety. Using these annotations, we fine-tuned a RuBERT-tinyDale (2021) filtration model and applied it to the deduplicated corpus. After filtration, the total number of tokens left for continued pre-training is 11 billion.
We observed that continued pre-training of a LLM can partially diminish the reasoning capabilities present in the original English-oriented model, significantly impacting overall performance. In our preliminary experiments, a model that underwent continued pre-training may demonstrate even worse performance on Russian benchmarks than the original English-oriented model. To alleviate this “catastrophic forgetting” in reasoning, we use the loss regularization with the KL penalty between the probability distribution of Vikhr and the reference English-oriented LLM:
@@
In practice, we implement the regularization using the SLERP interpolation of model losses Goddard et al. (2024).
@@
To speed up the process of continued pre-training, we use an optimized Flash attention implementation [^2] https://huggingface.co/docs/optimum/bettertransformer/tutorials/convert. As an optimization algorithm, we leverage AdamW, as it trades some memory efficiency in favor of robustness to the hyperparameter choice. The hyperparameters used for continued pre-training are presented in Table 3.
@@ 3.3 Instruction Tuning
Instruction tuning is an essential step in reaching high zero-shot performance with LLMs. It also allows to obtain a more natural communication with the model without complex prompting. Further fine-tuning techniques such as RLHF Ouyang et al. (2022) or DPO Rafailov et al. (2024), which require input from the assessors, are also crucial for such tasks as multicriteria alignment. However, the most significant performance gains are still achieved through instruction tuning Jha et al. (2023).
Previously, Gusev (2023) constructed an open-source set of instruction-output pairs for the Russian language (Saiga). The core Saiga dataset was created similar to Alpaca by querying ChatGPT (gpt-3.5-turbo) Taori et al. (2023). In this work, we extend this set by translating two English instruction datasets. First, we translated instructions for the FLAN model Wei et al. (2021) and generated answers in Russian using ChatGPT. Originally, FLAN instructions were constructed automatically from annotated datasets using templates to facilitate multitask and zero-shot capabilities of seq2seq models. Later, it was shown that this data also helps to improve decoder-only chat-oriented models as well. Second, we construct Veles [^3] https://huggingface.co/datasets/Vikhrmodels/Veles-2.5 by translating the English OpenHermes Teknium (2023) instruction dataset. Third, we incroporate without translation Nectar [^4] https://huggingface.co/datasets/berkeley-nest/Nectar Zhu et al. (2023) – an English instruction dataset. This ensures that Vikhr maintains strong performance in English as well.
Similar to the corpus for continued pre-training, we performed deduplication of the instruction set. Since the majority of the outputs were machine generated there are many low quality outputs. To mitigate this problem, we filtered out low quality pairs using a reward model trained on human data. For the reward model, we selected the multilingual-e5-large model Wang et al. (2024a). This model was particularly suitable for our needs due to its ability to handle multilingual data efficiently, ensuring that the classifier could accurately assess the quality of responses in both Russian and English. We trained the reward model on the answer preference dataset [^5] https://huggingface.co/datasets/Vikhrmodels/sbs, which was collected from human-written prompts and annotated using GPT-4. By applying this reward model, we filtered out low-quality instruction-output pairs, significantly enhancing the overall performance and reliability of our instruction datasets. The statistics of the Vikhr instruction datasets is presented in Table 4.
@@
Table 4: The statistics of the instruction datasets.
@@
(Ours) Vikhr-7B-instruct_0.2 true pt+sft 0.62 0.80 0.231 0.67 0.485
@@
Prompt Answer of base Mistral-0.2-Instruct Answer of Vikhr
Что такое чебурек?
[What does cheburek means?]
Чебурек - это классическое dish среднеазиатской и русской кухни, которое включает в себя тонкий lavash, завернутый в форме полумесяца или кольца. Он обычно наполнен meat (говядиной, бараниной или курицей) и cheese. История чебурека начинается с древних тюркских племен, которые готовили подобные dishes из мяса и теста. … Чебурек часто подается как горячее dish, жареное на сковороде или в духовке. Тесто обычно тонкое и хрустящее снаружи, а внутри мягкое и сочное. Мясо и сыр добавляют flavor depth и питательную ценность. Чебурек — это традиционное блюдо кавказской кухни, которое представляет собой большой, тонкий пирог из пресного теста с мясной начинкой, обжаренный во фритюре. Внутри чебурека обычно находится сочный фарш из баранины или говядины, приправленный луком, солью, перцем и другими специями. Благодаря обжарке в масле, тесто становится хрустящим снаружи, а начинка остается мягкой и сочной. История чебурека уходит корнями в древние времена, когда кочевники готовили его на раскалённых камнях или в больших котлах с горячим маслом.
Расскажи на русском языке что такое хачапури
[Tell in Russian what hatchapuri is]
Хачапури (хачпури, хачъапури) – это традиционное блюдо грузинской кухни. Это тип пицце или за Francia, выполненной в виде круглой или полукруглой лепёшки, покрытой верху layerм layerim – смесью тонко нарезанныхimo эplesses, лука .. Хачапури — это традиционное грузинское блюдо, представляющее собой лепешку с сыром внутри. Вот рецепт хачапури. Ингредиенты: - Для теста: - 3 стакана муки - 1 ч. л. соли - 1 ч. л. сахара - 100 мл теплой воды - Для смазывания: 1 ст. л. топленого масла …
Расскажи на русском языке как приготовить уху
[Tell in Russian how to cook a fish soup]
Уха (из Рыбы) – традиционное блюдо русской кухни, которое готовится predominantly из щуки или окуня. Рас**те процедуру приготовления ухи в домашних условиях: Выберите чистую рыбу. Очистить её от шерсти иICScoне всех мелких костей s… Уха — это традиционное русское блюдо, представляющее собой наваристый суп из рыбы, который варится на медленном огне. Вот базовый рецепт ухи, который можно адаптировать под свои вкусовые предпочтения. Ингредиенты: 1 кг свежей рыбы (можно использовать несколько видов, например, карп, щука, судак) 3-4 картофелины 1 крупная луковица 2-3 моркови …
Table 7: Comparison of texts generated by the base LLM Mistral-0.2-Instruct (on the left) with texts generated by our model Vikhr (on the right). Issues in the output are highlighted with red.